If a car of mass 400 kg reduces speed from 80 km/h to 34 km/h in 17s. Calculate the braking power required to give this change of speed.
1
Expert's answer
2020-11-23T10:27:26-0500

Let's first convert km/h to m/s:


"v_i=(80\ \dfrac{km}{h})\cdot (\dfrac{1000\ m}{1\ km})\cdot(\dfrac{1\ h}{3600\ s})=22.22\ \dfrac{m}{s},""v_f=(34\ \dfrac{km}{h})\cdot (\dfrac{1000\ m}{1\ km})\cdot(\dfrac{1\ h}{3600\ s})=9.44\ \dfrac{m}{s}."

We can find the braking power required to give this change of speed from the definition of the power:


"P=\dfrac{W}{t},"

here, "P" is the braking power, "W" is the work done by the braking force to change the speed from 80 km/h to 34 km/h, "t" is time.

We can find the work done by the braking force from the Work-Kinetic Energy Theorem:


"KE_f-KE_i=-W,""W=-(\dfrac{1}{2}mv_f^2-\dfrac{1}{2}mv_i^2)=-\dfrac{1}{2}m(v_f^2-v_i^2),""W=-\dfrac{1}{2}\cdot\ 400\ kg\ \cdot((9.44\ \dfrac{m}{s})^2-(22.22\ \dfrac{m}{s})^2)=80923\ J."

Finally, substituting "W" into the formula for the power we can calculate the braking power required to give this change of speed:


"P=\dfrac{W}{t}=\dfrac{80923\ J}{17\ s}=4760\ W=4.76\ kW."

Answer:

"P=4760\ W=4.76\ kW."